OpenStudy (anonymous):

i need someone to standby because i have a lot of questions please and thank you!

OpenStudy (anonymous):

given rectangle PQRS with PQ=QR, PQ=SP and SQ= 5 square root of 2, find RQ a) not enough info b)square root of 2 c)square root of 5 d)5 square root of 2

OpenStudy (anonymous):

Given that EFGH is an isosceles trapezoid, find x and y. A. x = 4, y= 2 B. x = 10, y= 46 C. x = 8,y = 2 D. x = 10, y= 44 http://media.education2020.com/evresources/2072371_5-9-2013_2-33-37_pm.jpg

OpenStudy (anonymous):

no need to worry pal u r at right place with many intellect persons to help u out.just pump in ur query.we will try our level best to sort them out.

OpenStudy (anonymous):

Which of the following statements is never true? A. A parallelogram is a quadrilateral. B. A trapezoid is a parallelogram. C. A square is a rhombus. D. A kite is a quadrilateral.

OpenStudy (anonymous):

and thank you so much

OpenStudy (anonymous):

Rhombus BOXY has angle measures as shown in the diagram. Find the measure of angle B A.30° B.60° C.100° D.120° http://media.education2020.com/evresources/2090848_f4e4ffac-f5f6-48fb-8a2b-94e88a38e67d.png

OpenStudy (anonymous):

oh i have to close this question because i can only ask one at a time

OpenStudy (anonymous):

you have to close this question and make other if not you will get a warning code of conduct

OpenStudy (anonymous):

i have already got one thank you

OpenStudy (anonymous):

lol sorry

OpenStudy (anonymous):

its fine im just taking an e2020 quiz and have to finish this class in 1 day now and im kinda worried i wont finish because you cant move on untill you pass the assesment

OpenStudy (anonymous):

answer to question 1. i got x=10 and y=46 degrees. since it is a isosceles trapezoid so 3y-4+y=180 property of quadrilateral for sum of opp side is 180 .so y=46 degree. then being an isosceles opp. skewed sides are equal,equate them and x=10

OpenStudy (anonymous):

thats number 2 :)

OpenStudy (anonymous):

question 2 i guess it option B as a parallelogram by its name has opposite sides parallel and equal unlike trapezoid having one set of opp side are equal but not parallel and other set parallel but not equal

OpenStudy (anonymous):

sorry for wrong question nos pal.arrange them pls

OpenStudy (anonymous):

for rhombus question,as its clearly given opp angles are equal. and sum of all angles of a quadrilateral is 360 degree so g+2g+g+2g=360 g=60 as angle B is 2g=2*60=120 degrees

OpenStudy (anonymous):

ty! i passed the quiz

OpenStudy (anonymous):

nice to hear these sweets words :) all the best for future